# Unsupervised Representation Learning for 3D Mesh Parameterization with Semantic and Visibility Objectives

### v2

@@ Abstract
Figure 1: 3D mesh parameterizations generated by our proposed semantic- (left) and visibility-aware (right) pipelines. Semantic-Aware (left - Sec. 3.3): To encourage semantically coherent UV charts that simplify texture editing, given an input 3D mesh (a), we design a partition-and-parameterize strategy: (b) compute a per-vertex semantic partition of the mesh, (c) learn a geometry-preserving UV parameterization (Sec. 3.2) independently for each semantic part to obtain per-part UV islands, and then aggregate and pack these islands into a unified UV atlas (insets). Visibility-Aware (right - Sec. 3.4): To encourage seamless UV mappings, our visibility-aware pipeline (d) takes an input 3D mesh, jointly (e) guides cutting-seam placement (red curves), extracts the corresponding boundary points in UV space (red dots), and (f) estimates a global geometry-preserving parameterization. As a result, the method steers cutting seams toward less-visible (more occluded) surface regions, resulting in more visually seamless UV maps.
@@ 1 Introduction
Texture mapping, also called UV/Mesh parameterization, converts a 3D mesh surface into a 2D image (atlas) and is a core step in digital 3D content production, enabling texture synthesis, painting, transfer, and high-quality rendering. While recent 3D generative methods produce high-quality textures for 3D mesh objects, they commonly rely on the heavy assumption that input 3D meshes are accompanied by the mesh parameterization process that requires both technical precision and artistic judgment. This assumption limits usability for content creators and designers. Moreover, industry surveys show that this process often accounts for a significant share of asset creation, creating a major bottleneck for 3D content creators. To address these issues and automate mesh parameterization, numerous methods have been proposed that mainly focus on geometric properties such as bijectivity Tutte (1963); Sheffer et al. (2007a), conformality (angle-preservation) Lévy et al. (2023), stretch (length-preservation) Sander et al. (2002), and equiareality (area-preservation) Zou et al. (2011). While these properties are necessary to achieve a high-quality parameterization, they are not sufficient for many downstream applications in content creation and texture synthesis. In particular, two perceptual criteria are often neglected: (1) semantic awareness, 2D UV charts should align with semantically meaningful surface 3D parts so that textures designed for a semantic region remain coherent and transferable across 3D shapes; and (2) visibility awareness, seams should be placed where they are unlikely to be observed under typical viewpoints and lighting, so that seam artifacts are less perceptible after texturing and rendering. To the best of our knowledge, only limited work has explored these properties individually: semantic-aware UV mapping Vermandere et al. (2024); Mukherjee et al. (2024) and visibility-aware UV learning Weill-Duflos et al. (2023); Sheffer & Hart (2002); Ray et al. (2010). However, none of the existing approaches address both objectives simultaneously. To overcome these shortcomings, we present an unsupervised differentiable framework that augments standard geometry-preserving UV learning with semantic- and visibility-aware objectives. For semantic-awareness, our pipeline (i) segments the mesh into semantic 3D parts, (ii) applies an unsupervised learned per-part UV-parameterization backbone, and (iii) aggregates per-part charts into a unified UV atlas. For visibility-awareness, we use ambient occlusion (AO) Zhukov et al. (1998) as an exposure proxy and back-propagate a soft differentiable AO-weighted seam objective designed to steer cutting seams toward occluded regions. The proposed pipeline has two stages: (i) a neural surface-parameterization backbone enforces geometry-preserving properties, (ii) semantic-aware and visibility-aware modules provide task-specific (e.g., texture painting) guidance. The main contributions of this study are: (i) We propose an unsupervised, differentiable framework that jointly estimates geometry-preserving UV parameterizations while optimizing semantic- and visibility-aware objectives. (ii) We introduce two novel perceptual objectives that directly connect UV parameterization to texture synthesis: a semantic-aware objective that aligns 2D UV charts with semantically meaningful 3D surface parts, and a visibility-aware objective that places seams in less-visible regions to minimize seam artifacts.
@@ 2 Related Work
Traditional UV Parameterization. These methods aim to flatten 3D surfaces to 2D while minimizing distortion and avoiding overlaps. Early conformal (angle-preserving) techniques such as LSCM Lévy et al. (2023) and ABF++ Sheffer et al. (2005) reduce angular distortion via global optimization or iterative angle adjustment. Many approaches also enforce bijectivity and bound distortion: bounded-distortion parameterization Sorkine et al. (2002) limits local stretch by cutting into patches, bijective free-boundary methods Smith & Schaefer (2015) avoid overlaps without fixing boundaries, and SLIM Rabinovich et al. (2017) scales these guarantees to large meshes. Optimizing cuts and chart segmentation also have been addressed: AutoCuts Poranne et al. (2017) and OptCuts Li et al. (2018) jointly optimize seams and UV parameterization to trade off distortion and chart count, while Seamster Sheffer & Hart (2002) and Seamless Liu et al. (2017) place seams to minimize visible texture artifacts by aligning them with features or low-visibility regions. Although classical algorithms provide the geometric foundation for UV mapping, they often treat these goals independently and rarely exploit higher-level objectives (e.g., semantic parts or visibility), which can limit final texture quality.
Learning-Based UV Parameterization. The existing methods for UV parameterization can be grouped by input representation: explicit (meshes and point clouds) Zhao et al. (2025); Zhang et al. (2024b); Srinivasan et al. (2024); Chen et al. (2022); Teimury et al. (2020); Oechsle et al. (2019); Knodt et al. (2023), implicit (NeRFs) Xu et al. (2024a); Xiang et al. (2021); Srinivasan et al. (2024); Guan et al. (2023); Das et al. (2022), and 3D-Gaussian-splatting-based methods Xu et al. (2024b). In this work, we focus on explicit (mesh) parameterization, which is widely used in industry. Mesh-based learning approaches Zhao et al. (2025); Srinivasan et al. (2024); Zhang et al. (2024a; b) usually train multi-layer perceptrons (MLPs) to learn UV mappings by optimizing three core criteria: (i) minimizing chart count, (ii) reducing distortion (preserving lengths, angles, and areas), and (iii) enforcing bijectivity via round-trip consistency. These MLPs are trained with carefully designed losses to satisfy those constraints. However, current learning-based methods often (i) lack semantic awareness, meaning that cutting seams are chosen only to reduce geometric distortion rather than to align with meaningful parts, and (ii) ignore visibility when placing seams, failing to steer cuts toward low-visibility regions and thus producing noticeable texture artifacts.
@@ 3 Methodology
Figure 2: An overview of the training process of the proposed semantic-aware UV parameterization method (Sec. 3.3), consisting of three main stage: (i) semantic 3D partitioning, where we compute a per-vertex semantic partition of the input mesh using shape diameter function (Appendix A.1); (ii) geometry-preserving UV learning, where we apply the base UV-parameterization backbone (Sec. 3.2) independently to each semantic part to obtain per-part UV islands; and (iii) UV atlas aggregation and packing aggregate and pack these islands into a unified UV atlas.
We propose a two-stage, unsupervised, end-to-end differentiable framework to automatically learn 3D mesh parameterizations with semantic and visibility objectives. The training consists of two stages. In stage (i), geometry-preservation mesh-parameterization learning (Sec. 3.2), an MLP-based network is trained with fully differentiable geometric objectives defined in the UV domain to produce an initial UV mapping that satisfies low-distortion requirements. In stage (ii), learning perceptual objectives (Sec. 3.3 and Sec. 3.4), we introduce two perceptual objectives: semantic awareness and visibility awareness. By semantic awareness, we mean that the UV charts produced by stage (i) should align with semantically meaningful surface 3D parts. To this end, we adopt a partition-and-parameterize strategy based on the Shape Diameter Function Shapira et al. (2008): we first partition the input mesh into semantic 3D parts, learn a UV parameterization for each segmented part with the same MLP-based architecture used in stage (i), and finally aggregate the per-part UV charts into a single unified UV atlas. By visibility awareness, we mean that cutting seams should be placed on 3D mesh surface regions that are unlikely to be seen under typical viewpoints and lighting, making seam artifacts less perceptible after texturing. To this end, we propose a visibility-aware objective that uses ambient occlusion as a proxy for viewer exposure Zhukov et al. (1998), encouraging seam placement in regions of low exposure. An overview of the training procedure for semantic awareness and visibility awareness pipelines is shown in Fig. 2 and Algorithm 1, respectively.
@@ 3.1 Problem Statement and Preliminaries
Given a 3D mesh $M=(V,F)$ with vertex positions $V\in\mathbb{R}^{N_{v}\times 3}$ (where $N_{v}$ is the number of vertices), triangular faces $F\in\mathbb{R}^{N_{f}\times 3}$ (where $N_{f}$ is the number of faces), and per-vertex normals $N\in\mathbb{R}^{N_{v}\times 3}$, our goal is to learn a pointwise UV parameterization $u_{\theta}:V\rightarrow\mathbb{R}^{2}$ that maps each vertex in $V$ to a 2D coordinate, producing UV coordinates $Q\in\mathbb{R}^{N_{v}\times 2}$ with row-wise correspondence between $V$ and $Q$. We require the learned parameterization to satisfy the following properties during the learning process: (i) geometry preservation, i.e., the parameterization should be low-distortion and ideally preserve classical geometric properties such as (a) isometry (maintaining original edge lengths), (b) equiareal (preserving triangle/facet areas), and (c) conformality (preserving angles between edges); (ii) UV-space efficiency and bijectivity, the mapping should (a) efficiently utilize the available UV domain, (b) avoid overlapping projections of distinct 3D points, and (c) be one-to-one so that each surface point maps to a unique UV location; (iii) high-quality cutting seams, the parameterization should place seams in visually inconspicuous regions Weill-Duflos et al. (2023); Ray et al. (2010); Sheffer & Hart (2002); and (iv) semantic awareness, the UV mapping should semantically align 3D parts across different 3D mesh models to consistent regions in UV space to enable cross-shape correspondence Vermandere et al. (2024); Mukherjee et al. (2024); Cohen-Bar et al. (2025).
@@ 3.2 Base Neural Architecture for UV-Parameterization Learning
To emulate the sequence of operations an artist would typically perform when constructing a UV atlas for a 3D mesh, such as stretching a canvas, placing it over the 3D surface, adding surface cuts where needed, and finally flattening it back to the plane, we build upon the idea of a deforming a 2D grid Yang et al. (2018) by leveraging bi-directional cycle mapping backbone, proposed in Zhang et al. (2024a; b); Srinivasan et al. (2024); Zhao et al. (2025). We emphasize that this stage adopts an existing previously proposed low-distortion parameterization backbone. We do not claim novelty for the backbone itself. Instead, our primary contribution lies in the design and integration semantic-aware and visibility-aware objectives as downstream tasks built on top of this established geometric foundation. Intuitively, the bi-directional cycle mapping enforces that projecting a 3D surface patch into 2D UV coordinates and re-projecting those UVs back onto the 3D mesh yields the original data. Therefore, during the learning process, any seams, local stretches, distortions, or semantic misalignments show up as reconstruction errors. Then, by minimizing these errors, the model is guided to produce an approximately invertible and coherent parameterization that minimizes distortion. To build such a mapping, inspired by recent advances in global parameterization networks Zhao et al. (2025); Srinivasan et al. (2024), our backbone is composed of four geometrically interpretable subnetworks: (i) deforming network, which starts from a regular 2D grid and adaptively deforms it into candidate UV coordinates; (ii) wrapping network, which maps those 2D UV candidates to the 3D surface by smoothly bending the deformed grid to conform to the input mesh, bridging the 2D and 3D domains; (iii) cutting network, which operates on the wrapped 3D surface to open the closed geometry by placing seams that reduce distortion for subsequent flattening; and (iv) unwrapping network, which projects the resulting open 3D surface back to 2D while preserving smoothness and geometric coherence. In this design, Each subnetwork is a pointwise MLP applied to mesh vertices, and the whole architecture is trained end-to-end in an unsupervised fashion. The four subnetworks form two complementary branches enforcing 2D–3D–2D and 3D–2D–3D cycles for ensuring consistency across 2D and 3D domains. Coupled with differentiable losses that encourage bijectivity and minimize distortion, this backbone produces candidate UV parameterizations for the next stages. Further details on the loss design and the interaction of the four sub-networks within the two-branch bi-directional cycle are given in Appendix A.2.
@@ 3.3 Semantic-Aware UV Parameterization
The UV parameterization, obtained from the bi-directional cycle mapping, is used in the next step, semantic-aware UV parameterization. To encourage semantically coherent UV charts that simplify texture editing, transfer, and cross-object correspondence, we design a partition-and-parameterize pipeline (illustrated in Fig. 2). Our semantic-aware parameterization has three stages: (i) compute a per-vertex semantic partition of the input mesh using the shape diameter function (ShDF) Shapira et al. (2008); (ii) apply the base UV-parameterization backbone from Sec. 3.2 independently to each semantic 3D part to obtain per-part UV islands; and (iii) aggregate and pack these islands into a unified UV atlas. We now describe each stage in detail.
@@
3D Partitioning using Shape Diameter Function. Given an input mesh $\mathcal{M}=(V,F)$, we compute a semantic partition $S:V\to\{1,\dots,K\}$ using the shape diameter function (ShDF) Shapira et al. (2008) as our primary signal. Intuitively, the ShDF maps each surface sample to a scalar that estimates the local object thickness (the diameter of the object in the vicinity of that sample). Similar ShDF values typically indicate coherent semantic parts (e.g., limbs, body, handles), making the ShDF a practical cue for part-level partitioning of 3D shapes. Specifically, our 3D partitioning strategy consists of the following steps (See Appendix. A.1 for the detailed explanation of each step): (i) compute local ShDF, we compute a local thickness (ShDF) per face via ray casting inside a cone about the inward normal at each surface point and smoothing the samples for spatial coherence; (ii) fit GMM, to obtain soft per-face class likelihoods, we use a 1-D Gaussian mixture model (GMM) to fit $K$ (the deisred number of semantic components) Gaussians to the per-face thicknesses; (iii) boundary smoothness cost, to bias the partition toward spatially coherent regions, we then combine these likelihoods with a geometry-aware pairwise smoothness penalty (e.g., a dihedral-angle based boundary cost) on adjacent faces on the geometry to encourage coherent regions and penalize label changes across sharp edges. (iv) initial labeling, we then assign each face to the GMM component with maximum probability (minimum negative log-likelihood). This produces a purely data-driven segmentation that is generally noisy but captures the major mode structure of the thickness field. (v) refine with graph cuts, we then refine the initial assignment by minimizing a global energy (see Eq. 10 in Appendix A.1) via iterative alpha-expansion, solving each expansion by an optimal min-cut. Repeating this process across labels and iterating a few times yields stable and low-energy 3D partitions. (vi) Relabel connected components, after refinement, a single label index can correspond to multiple disconnected components. To solve this issue, we relabel each connected component of faces so that every final label corresponds to a single connected sub-mesh. This guarantees that subsequent per-part processing operates on contiguous regions. (vii) Post-processing, to produce clean 3D segmentation results, we apply several simple post-processing heuristics to remove spurious tiny components that would complicate per-part UV parameterization. The output is then a set of connected semantic sub-meshes $\{\mathcal{M}_{k}=(V_{k},F_{k})\}_{k=1}^{K}$ suitable for the per-part UV parameterization stage described in Sec.3.2. This high-level pipeline is simple, robust, and fast. Full algorithmic and implementation details are provided in Appendix A.1 for reproducibility. We also evaluate a modern zero-shot segmentation baseline, SAMesh Tang et al. (2024), as an alternative partitioner and report comparative results in our ablation study.
Algorithm 1 Visibility-Aware UV Parameterization Training
@@
2:  # Step 1: Learning Mesh Parameterization
@@
Atlas Aggregation and Packing. After optimizing all parts, we obtain per-part UV islands $\{Q_{k}\subset[0,1]^{2}\}_{k=1}^{K}$ (each normalized to the unit square) corrsponding to each 3D segmented part. These islands are merged into a single atlas by an atlas aggregator. For clarity and reproducibility our current aggregator is intentionally simple and deterministic: the final UV sheet is a unit square subdivided into a $G\times G$ regular grid, where $G\;=\;\big\lceil\sqrt{K}\,\big\rceil$. We then assign each part $k$ a unique grid cell indexed by row and column $(r_{k},c_{k})$ in row-major order. Each normalized island $Q_{k}$ is placed into its assigned cell via a uniform similarity transform (scale $s$ and translation $t_{r_{k},c_{k}}$):

### v3

# arXiv:2509.25094v3[cs.GR] 27 Feb 2026

@@ Abstract
Figure 1: 3D mesh parameterizations generated by our semantic-aware (left) and visibility-aware (right) pipelines. Semantic-Aware (left, Sec. 3.3): To encourage semantically coherent UV charts for easier texture editing, given an input mesh (a), we design a partition-and-parameterize strategy, (b) compute a per-vertex semantic partition, (c) learn a geometry-preserving UV parameterization (Sec. 3.2) independently for each part to obtain UV islands, and finally aggregate and pack these islands into a unified UV atlas (insets). Visibility-Aware (right, Sec. 3.4): To encourage seamless UV mappings, the pipeline (d) takes an input 3D mesh, jointly (e) guides cutting-seam placement (red curves), extracts corresponding UV boundary points (red dots), and (f) estimates a global geometry-preserving parameterization. This steers cutting seams toward less-visible (more occluded) surface regions, producing more visually seamless UV maps.
@@ 1 Introduction
Texture mapping, also called UV/Mesh parameterization, converts a 3D mesh surface into a 2D image and is a core step in digital 3D content production, enabling texture synthesis, painting, transfer, and high-quality rendering. While recent 3D generative methods produce high-quality textures for 3D mesh objects, they commonly rely on the heavy assumption that input 3D meshes are accompanied by the mesh parameterization process that requires technical precision and artistic judgment. This assumption limits usability for content creators and designers. Moreover, industry surveys show that this process often accounts for a significant share of asset creation, creating a major bottleneck for 3D content creators. To address these issues, numerous methods have been proposed that mainly focus on geometric properties such as bijectivity (Tutte, 1963; Sheffer et al., 2007a), conformality (angle-preservation) (Lévy et al., 2023), stretch (length-preservation) (Sander et al., 2002), and equiareality (area-preservation) (Zou et al., 2011). While these properties are necessary to achieve a high-quality parameterization, they are not sufficient for many downstream applications in content creation and texture synthesis. In particular, two perceptual criteria are often neglected: (1) semantic awareness, 2D UV charts should align with semantically meaningful surface 3D parts so that textures designed for a semantic region remain coherent and transferable across 3D shapes; and (2) visibility awareness, seams should be placed where they are unlikely to be observed under typical viewpoints and lighting, so that seam artifacts are less perceptible after texturing and rendering. To the best of our knowledge, only limited work has explored these properties individually: semantic-aware UV mapping (Vermandere et al., 2024; Mukherjee et al., 2024) and visibility-aware UV learning (Weill-Duflos et al., 2023; Sheffer & Hart, 2002; Ray et al., 2010). However, none of the existing approaches address both objectives simultaneously. To overcome these shortcomings, we present an unsupervised differentiable framework that augments standard geometry-preserving UV learning with semantic- and visibility-aware objectives. For semantic-awareness, our pipeline (i) segments the mesh into semantic 3D parts, (ii) applies an unsupervised learned per-part UV-parameterization backbone, and (iii) aggregates per-part charts into a unified UV atlas. For visibility-awareness, we use ambient occlusion (AO) (Zhukov et al., 1998) as an exposure proxy and back-propagate a soft differentiable AO-weighted seam objective designed to steer cutting seams toward occluded regions. The proposed pipeline has two stages: (i) a neural surface-parameterization backbone enforces geometry-preserving properties, (ii) semantic-aware and visibility-aware modules provide task-specific (e.g., texture painting) guidance. The main contributions of this study are: (i) We propose an unsupervised, differentiable framework that jointly estimates geometry-preserving UV parameterizations while optimizing semantic- and visibility-aware objectives. (ii) We introduce two novel perceptual objectives that directly connect UV parameterization to texture synthesis: a semantic-aware objective that aligns 2D UV charts with semantically meaningful 3D surface parts, and a visibility-aware objective that places seams in less-visible regions to minimize seam artifacts.
@@ 2 Related Work
Traditional UV Parameterization. These methods aim to flatten 3D surfaces to 2D while minimizing distortion and avoiding overlaps. Early conformal (angle-preserving) techniques such as LSCM (Lévy et al., 2023) and ABF++ (Sheffer et al., 2005) reduce angular distortion via global optimization or iterative angle adjustment. Many approaches also enforce bijectivity and bound distortion: bounded-distortion parameterization (Sorkine et al., 2002) limits local stretch by cutting into patches, bijective free-boundary methods (Smith & Schaefer, 2015) avoid overlaps without fixing boundaries, and SLIM (Rabinovich et al., 2017) scales these guarantees to large meshes. Optimizing cuts and chart segmentation also have been addressed: AutoCuts (Poranne et al., 2017) and OptCuts (Li et al., 2018) jointly optimize seams and UV parameterization to trade off distortion and chart count, while Seamster (Sheffer & Hart, 2002) and Seamless (Liu et al., 2017) place seams to minimize visible texture artifacts by aligning them with features or low-visibility regions. Although classical algorithms provide the geometric foundation for UV mapping, they often treat these goals independently and rarely exploit higher-level objectives (e.g., semantic parts or visibility), which can limit final texture quality.
Learning-Based UV Parameterization. The existing methods for UV parameterization can be grouped by input representation: explicit (meshes and point clouds) (Zhao et al., 2025; Zhang et al., 2024b; Srinivasan et al., 2024; Chen et al., 2022; Teimury et al., 2020; Oechsle et al., 2019; Knodt et al., 2023), implicit (NeRFs) (Xu et al., 2024a; Xiang et al., 2021; Srinivasan et al., 2024; Guan et al., 2023; Das et al., 2022), and 3D-Gaussian-splatting-based methods (Xu et al., 2024b). In this work, we focus on explicit (mesh) parameterization, which is widely used in industry. Mesh-based learning approaches (Zhao et al., 2025; Srinivasan et al., 2024; Zhang et al., 2024a; b) usually train multi-layer perceptrons (MLPs) to learn UV mappings by optimizing three core criteria: (i) minimizing chart count, (ii) reducing distortion (preserving lengths, angles, and areas), and (iii) enforcing bijectivity via round-trip consistency. These MLPs are trained with carefully designed losses to satisfy those constraints. However, current learning-based methods often (i) lack semantic awareness, meaning that cutting seams are chosen only to reduce geometric distortion rather than to align with meaningful parts, and (ii) ignore visibility when placing seams, failing to steer cuts toward low-visibility regions and thus producing noticeable texture artifacts.
@@ 3 Methodology
Figure 2: An overview of the training of the proposed semantic-aware UV parameterization method (Sec. 3.3), consisting of three stages: (i) semantic 3D partitioning, computes a per-vertex semantic partition of the input mesh using shape diameter function (Appendix A.1); (ii) geometry-preserving UV learning, applies the base UV-parameterization backbone (Sec. 3.2) independently to each semantic part to obtain per-part UV islands; and (iii) UV atlas aggregation and packing, aggregates and packs these islands into a unified UV atlas.
We propose a two-stage, unsupervised framework to automatically learn 3D mesh parameterizations with semantic and visibility objectives. The training consists of two stages. In stage (i), geometry-preservation mesh-parameterization learning (Sec. 3.2), an MLP-based network is trained with fully differentiable geometric objectives defined in the UV domain to produce an initial UV mapping that satisfies low-distortion requirements. In stage (ii), learning perceptual objectives (Sec. 3.3 and Sec. 3.4), we introduce two perceptual objectives: semantic awareness and visibility awareness. By semantic awareness, we mean that the UV charts produced by stage (i) should align with semantically meaningful surface 3D parts. To this end, we adopt a partition-and-parameterize strategy based on the Shape Diameter Function (Shapira et al., 2008): we first partition the input mesh into semantic 3D parts, learn a UV parameterization for each segmented part with the same MLP-based architecture used in stage (i), and finally aggregate the per-part UV charts into a single unified UV atlas. By visibility awareness, we mean that cutting seams should be placed on 3D mesh surface regions that are unlikely to be seen under typical viewpoints and lighting, making seam artifacts less perceptible after texturing. To this end, we propose a visibility-aware objective that uses ambient occlusion as a proxy for viewer exposure (Zhukov et al., 1998), encouraging seam placement in regions of low exposure. An overview of the training procedure for semantic awareness and visibility awareness pipelines is shown in Fig. 2 and Algorithm 1, respectively.
@@ 3.1 Problem Statement and Preliminaries
Given a 3D mesh $M=(V,F)$ with vertex positions $V\in\mathbb{R}^{N_{v}\times 3}$ (where $N_{v}$ is the number of vertices), triangular faces $F\in\mathbb{R}^{N_{f}\times 3}$ (where $N_{f}$ is the number of faces), and per-vertex normals $N\in\mathbb{R}^{N_{v}\times 3}$, our goal is to learn a pointwise UV parameterization $u_{\theta}:V\rightarrow\mathbb{R}^{2}$ that maps each vertex in $V$ to a 2D coordinate, producing UV coordinates $Q\in\mathbb{R}^{N_{v}\times 2}$ with row-wise correspondence between $V$ and $Q$. We require the learned parameterization to satisfy the following properties during the learning process: (i) geometry preservation, i.e., the parameterization should be low-distortion and ideally preserve classical geometric properties such as (a) isometry (maintaining original edge lengths), (b) equiareal (preserving triangle/facet areas), and (c) conformality (preserving angles between edges); (ii) UV-space efficiency and bijectivity, the mapping should (a) efficiently utilize the available UV domain, (b) avoid overlapping projections of distinct 3D points, and (c) be one-to-one so that each surface point maps to a unique UV location; (iii) high-quality cutting seams, the parameterization should place seams in visually inconspicuous regions (Weill-Duflos et al., 2023; Ray et al., 2010; Sheffer & Hart, 2002); and (iv) semantic awareness, the UV mapping should semantically align 3D parts across different 3D mesh models to consistent regions in UV space to enable cross-shape correspondence (Vermandere et al., 2024; Mukherjee et al., 2024; Cohen-Bar et al., 2025).
@@ 3.2 Base Neural Architecture for UV-Parameterization Learning
To emulate the sequence of operations an artist would typically perform when constructing a UV atlas for a 3D mesh, such as stretching a canvas, placing it over the 3D surface, adding surface cuts where needed, and finally flattening it back to the plane, we build upon the idea of a deforming a 2D grid (Yang et al., 2018) by leveraging bi-directional cycle mapping backbone, proposed in (Zhang et al., 2024a; b; Srinivasan et al., 2024; Zhao et al., 2025). We emphasize that this stage adopts an existing previously proposed low-distortion parameterization backbone. We do not claim novelty for the backbone itself. Instead, our primary contribution lies in the design and integration semantic-aware and visibility-aware objectives as downstream tasks built on top of this established geometric foundation. Intuitively, the bi-directional cycle mapping enforces that projecting a 3D surface patch into 2D UV coordinates and re-projecting those UVs back onto the 3D mesh yields the original data. Therefore, during the learning process, any seams, distortions, or semantic misalignments show up as reconstruction errors. Then, by minimizing these errors, the model is guided to produce an approximately invertible and coherent parameterization that minimizes distortion. To build such a mapping, inspired by recent advances in global parameterization networks (Zhao et al., 2025; Srinivasan et al., 2024), our backbone is composed of four geometrically interpretable subnetworks: (i) deforming network, which starts from a regular 2D grid and adaptively deforms it into candidate UV coordinates; (ii) wrapping network, which maps those 2D UV candidates to the 3D surface by smoothly bending the deformed grid to conform to the input mesh, bridging the 2D and 3D domains; (iii) cutting network, which operates on the wrapped 3D surface to open the closed geometry by placing seams that reduce distortion for subsequent flattening; and (iv) unwrapping network, which projects the resulting open 3D surface back to 2D while preserving smoothness and geometric coherence. In this design, Each subnetwork is a pointwise MLP applied to mesh vertices, and the whole architecture is trained end-to-end in an unsupervised fashion. The four subnetworks form two complementary branches enforcing 2D–3D–2D and 3D–2D–3D cycles for ensuring consistency across 2D and 3D domains. Coupled with differentiable losses that encourage bijectivity and minimize distortion, this backbone produces candidate UV parameterizations for the next stages. Further details on the loss design and the interaction of the four sub-networks within the two-branch bi-directional cycle are given in Appendix A.2.
@@ 3.3 Semantic-Aware UV Parameterization
The UV parameterization, obtained from the bi-directional cycle mapping, is used in the next step, semantic-aware UV parameterization. To encourage semantically coherent UV charts that simplify texture editing, transfer, and cross-object correspondence, we design a partition-and-parameterize pipeline (illustrated in Fig. 2). Our semantic-aware parameterization has three stages: (i) compute a per-vertex semantic partition of the input mesh using the shape diameter function (ShDF) (Shapira et al., 2008); (ii) apply the base UV-parameterization backbone from Sec. 3.2 independently to each semantic 3D part to obtain per-part UV islands; and (iii) aggregate and pack these islands into a unified UV atlas. We now describe each stage in detail.
@@
3D Partitioning using Shape Diameter Function. Given an input mesh $\mathcal{M}=(V,F)$, we compute a semantic partition $S:V\to\{1,\dots,K\}$ using the shape diameter function (ShDF) (Shapira et al., 2008) as our primary signal. Intuitively, the ShDF maps each surface sample to a scalar that estimates the local object thickness (the diameter of the object in the vicinity of that sample). Similar ShDF values typically indicate coherent semantic parts (e.g., limbs, body, handles), making the ShDF a practical cue for part-level partitioning of 3D shapes. Specifically, our 3D partitioning strategy consists of the following steps (See Appendix. A.1 for the detailed explanation of each step): (i) compute local ShDF, we compute a local thickness (ShDF) per face via ray casting inside a cone about the inward normal at each surface point and smoothing the samples for spatial coherence; (ii) fit GMM, to obtain soft per-face class likelihoods, we use a 1-D Gaussian mixture model (GMM) to fit $K$ (the deisred number of semantic components) Gaussians to the per-face thicknesses; (iii) boundary smoothness cost, to bias the partition toward spatially coherent regions, we then combine these likelihoods with a geometry-aware pairwise smoothness penalty (e.g., a dihedral-angle based boundary cost) on adjacent faces on the geometry to encourage coherent regions and penalize label changes across sharp edges. (iv) initial labeling, we then assign each face to the GMM component with maximum probability (minimum negative log-likelihood). This produces a purely data-driven segmentation that is generally noisy but captures the major mode structure of the thickness field. (v) refine with graph cuts, we then refine the initial assignment by minimizing a global energy (see Eq. 10 in Appendix A.1) via iterative alpha-expansion, solving each expansion by an optimal min-cut. Repeating this process across labels and iterating a few times yields stable and low-energy 3D partitions. (vi) Relabel connected components, after refinement, a single label index can correspond to multiple disconnected components. To solve this issue, we relabel each connected component of faces so that every final label corresponds to a single connected sub-mesh. This guarantees that subsequent per-part processing operates on contiguous regions. (vii) Post-processing, to produce clean 3D segmentation results, we apply several simple post-processing heuristics to remove spurious tiny components that would complicate per-part UV parameterization. The output is then a set of connected semantic sub-meshes $\{\mathcal{M}_{k}=(V_{k},F_{k})\}_{k=1}^{K}$ suitable for the per-part UV parameterization stage described in Sec.3.2. This high-level pipeline is simple, robust, and fast. Full algorithmic and implementation details are provided in Appendix A.1 for reproducibility. We also evaluate a modern zero-shot segmentation baseline, SAMesh (Tang et al., 2024), as an alternative partitioner and report comparative results in our ablation study.
Algorithm 1 Visibility-Aware UV Learning
@@
2:  # Step 1: Learning UV map
@@
Atlas Aggregation and Packing. After optimizing all parts, we obtain per-part UV islands $\{Q_{k}\subset[0,1]^{2}\}_{k=1}^{K}$ (each normalized to the unit square) corrsponding to each 3D segmented part.
These islands are merged into a single atlas by an atlas aggregator. For clarity and reproducibility our current aggregator is intentionally simple and deterministic: the final UV sheet is a unit square subdivided into a $G\times G$ regular grid, where $G\;=\;\big\lceil\sqrt{K}\,\big\rceil$. We then assign each part $k$ a unique grid cell indexed by row and column $(r_{k},c_{k})$ in row-major order. Each normalized island $Q_{k}$ is placed into its assigned cell via a uniform similarity transform (scale $s$ and translation $t_{r_{k},c_{k}}$):